OpenStudy (melissa_something):

Choose the equation below that represents the line passing through the point (-3, -1) with a slope of 4

OpenStudy (anonymous):

Use this formula: y-y1=m(x-x1)

OpenStudy (melissa_something):

y + 1 = 4(x + 3) ? Idk if its positive because (-3,-1) are both negative -.-

OpenStudy (anonymous):

Yes, y + 1 = 4(x + 3) is right.

OpenStudy (melissa_something):

omg thank you!!!(:

OpenStudy (anonymous):

I'm not sure what the choices are, but you probably need to distribute the 4 and solve for y.

OpenStudy (anonymous):

No prob. Glad I can help.

OpenStudy (melissa_something):

Its y = 4x - 1 y = 4x - 3 y = x - 3 y = x - 1

OpenStudy (anonymous):

I got y=4x+11 ...

OpenStudy (melissa_something):

Oh wait major fail wrong answeres y - 1 = 4(x - 3) y + 1 = 4(x + 3) y - 3 = 4(x + 1) y + 3 = 4(x - 1) my bad lol

OpenStudy (melissa_something):

But would it be y + 1 = 4(x + 3) because 2 negatives make a positive ?